### v1

@@ 3.2 Time series data
There already exists a variety of techniques to tackle time series problems. For many examples, one can often get reasonable predictions by using relatively simple models such as auto-regressive moving average models and their many variants [129]. For time series data with additional challenges (such as the ones mentioned at the start of this section), the current state-of-the-art is the long short-term memory (LSTM) ML model [130], which is able to capture some properties, such as seasonality and highly correlated variables, much better than the more traditional models. However, a recent QML technique that has also started to gain attention is quantum reservoir computing (QRC) [131]. Reservoir computing is a predecessor to recurrent neural networks (RNNs), where the hidden weights can be treated as a reservoir that projects the data into a higher-dimensional subspace, which can then be fed into a simple linear model; quantum analogues exist [132]. The reservoir (similar to the RNN) receives the time series data sequentially, and hence the dynamics of the reservoir are dependent on the previous time step. This reservoir can have many different forms, ranging from classical architectures like the echo state network [133] to physical implementations [134, 135]. The advantage of reservoir computing versus RNNs is that since the majority of the randomly initialized parameters of this reservoir is fixed from the start, the computational cost of training them is much lower. On the other hand, this randomness typically requires reservoirs to have many more parameters to achieve a comparable accuracy. Still, the use of a quantum reservoir (for example, by generating a random quantum circuit) can decrease the number of parameters required for the model to perform well.
@@ 3.2.1 EHRs
Furthermore, wearables represent another modality that is growing rapidly and which is starting to be integrated with EHRs [137]. These include chemical sensors that collect signals with electrochemical and optical methods [138] as well as fitness trackers and other devices which monitor digital biomarkers [139]. They may also include quantum sensors that exploit quantum mechanical properties in order to achieve superior sensitivities, such as nitrogen–vacancy-centre-based magnetometry at the level of single neurons [140]. By their nature, these wearables typically generate time series data and could thus benefit from such quantum (machine learning) methods.
@@ 3.2.2 Omics
Omics time series data is particularly valuable for biomarker discovery, as it allows one to observe temporal patterns and gene-gene/protein-protein interactions. Examples includes disease progression/sub-typing [141, 142] and patient response to treatments [143, 144, 145, 146].
Non-invasive approaches to use liquid biopsies allow one to capture a mixture of the omics profile of cancer cells in a patient via cell-free DNA (cfDNA). Recently, studies have shown application of cfDNA to identify drug resistance mechanisms [147]. Moreover, in the last few years several studies have shown that cfDNA fragmentomics’ characteristics differ in normal and diseased individuals without the need to distinguish the source of the cfDNA fragments, which makes it a promising novel biomarker in particular for early cancer detection [148, 149]. Indeed, ML algorithms have been applied to such tasks with promising results to identify late-stage cancer [148] but still fail to detect early signs in the development of the disease. Another recent application of omics time series is in the case of epilepsy post-surgical seizure recurrence [141] where omics data were used to subtype patients based on their treatment responses.
In general, classical ML often cannot be easily applied to omics-based time series data due to issues such as small sample numbers or missing data. Therefore, data imputation is required, and current imputation techniques often do not incorporate time-dependence, thereby leaving out critical information during imputation. Recent work has demonstrated that graph networks that model topology across samples can be hugely beneficial in such cases [150]. An example is the use of dynamic Bayesian networks for analysis of time series microbiome data [151], which has been demonstrated to be a critical biomarker for many diseases states [152, 153]. Still, these topological methods quickly become computationally infeasible with increasingly high-dimensional and complicated data sets, which is where QML and quantum topological data analysis (QTDA) [154] approaches may be able to help, providing an orthogonal point of view and insights. Indeed, they could be used to differentiate with high accuracy between healthy individuals and cancer patients in early stages of diseases, leading to more targeted and effective treatments.
@@ 3.2.3 Medical images
QRC techniques are relevant techniques for application to these problems. However, the quantum computers would have to be larger than the current generations or the initial image data would have to be downscaled with classical pre-processing techniques, such as passing the image through a residual network [155] before sending it through the QRC model. Further investigation is required to understand whether the loss of information during the classical pre-processing would negate the potential advantages of using QRC and related quantum methods instead of classical techniques.
@@ 3.3 Erroneous data
Noise and errors often appear in the context of quantum computing due to the fragility of the quantum systems and their inherent tendency to decohere. A variety of approaches is being researched in order to handle, or even leverage, such noise [156, 157, 12, 158, 159] and progress towards quantum error correction [160], arguably the holy grail of quantum computing. The focus of this discussion, however, is on handling noisy datasets where the labels are incorrect or missing, which represents a major challenge for classical algorithms [161, 162, 163]. Such misannotations are also of concern when data is synthetically generated [164], an area that is also being explored with QGANs and related methods [165].
@@ 3.3.1 EHRs
EHR data has been notoriously difficult to analyze due to factors such as errors, lack of interoperability, record gaps, and unusually large or small data sizes. While leveraging such data may also allow better generalization behavior through classical models [81], QML has shown promise in dealing with such difficult datasets even more effectively. For example, the usual presence of quantum noise may actually lead to a greater degree of global robustness and thus fairer models [166]. Moreover, quantum transfer learning may be beneficial when the (real-world) dataset of interest is particularly low- or high-dimensional [167]. Quantum support vector machines have been applied to EHRs and have, for small datasets, been shown to be competitive with classical approaches in classifying ischemic heart disease [168] and predicting the persistence of rheumatoid arthritis patients [169]. Small datasets, for instance for rare-disease or other very restricted cohorts, may in fact lead to some of the earliest quantum advantages, given that quantum generative and other QML models are particularly suitable for such a setting [28, 170]. Finally, the process of imputation itself may be enhanced, leveraging quantum determinantal point processes [171].
@@ 4 Open research challenges
In the space of biomarker discovery, the loading of data is one key challenge. As described before, biomarker discovery often necessitates extracting insights from large classical datasets. Despite progress with QRAMs, breakthroughs in this area will unlock many more applications. Related to that, quantum algorithms remain an intensely researched area. New ones are being continually discovered; for instance, a study recently claimed an exponential advantage in pathfinding for graphs [178]. Furthermore, quantum federated learning represents another novel area [179, 180, 181, 182]. The idea is to enhance federated learning models where different entities train a model while keeping their data decentralized. Given the growing importance federated learning has for healthcare and medicine, from allowing multi-institutional collaborations without patient data sharing [183] to enabling digital health [184] to enhancing the internet of medical things [185], this is a clear innovation opportunity.
In the area of biomarker validation, a critical aspect concerns benchmarks. The challenge with demonstrating value from quantum computing is that the (classical) goal posts are often not clearly defined and may shift (particularly in response to, or even inspired by, quantum algorithm advances). This makes it challenging to justify the efforts required to develop and use quantum algorithm-based tools. Connected with that, generalization is a key issue due to individual biomarker variability resulting from genetic and lifestyle factors [186]. As a result, more granular (quantum) models will have to be developed, which might again require more (classical) data and novel quantum algorithms.
@@ 5 Conclusion
Clearly, there are also still discrepancies with regard to the access to advanced forms of computing, such as quantum computing, across the world [187]. While quantum computing, in particular, has been rapidly made more widely available, it is likely to remain relatively expensive for some time to come; today, typical costs are thousands of US dollars per QPU hour. Thus, democratization of such transformative technologies is an important issue that also needs to be considered in order to realize their full impact across the world [188].

### v2

@@ 3.2 Time series data
There already exists a variety of techniques to tackle time series problems. For many examples, one can often get reasonable predictions by using relatively simple models such as auto-regressive moving average models and their many variants. For time series data with additional challenges (such as the ones mentioned at the start of this section), the current state-of-the-art is arguably the long short-term memory (LSTM) ML model, which is able to capture some properties, such as seasonality and highly correlated variables, much better than the more traditional models. However, a recent QML technique that has also started to gain attention is quantum reservoir computing (QRC) [69]. Reservoir computing is a predecessor to recurrent neural networks (RNNs), where the hidden weights can be treated as a reservoir that projects the data into a higher-dimensional subspace, which can then be fed into a simple linear model; quantum analogues exist [70]. The reservoir (similar to the RNN) receives the time series data sequentially, and hence the dynamics of the reservoir are dependent on the previous time step. This reservoir can have many different forms, ranging from classical architectures like the echo state networkto physical implementations. The advantage of reservoir computing versus RNNs is that since the majority of the randomly initialized parameters of this reservoir is fixed from the start, the computational cost of training them is much lower. On the other hand, this randomness typically requires reservoirs to have many more parameters to achieve a comparable accuracy. Still, the use of a quantum reservoir (for example, by generating a random quantum circuit) can decrease the number of parameters required for the model to perform well.
@@ 3.2.1 EHRs
Furthermore, wearables represent another modality that is growing rapidly and which is starting to be integrated with EHRs. These include chemical sensors that collect signals with electrochemical and optical methodsas well as fitness trackers and other devices which monitor digital biomarkers. They may also include quantum sensors that exploit quantum mechanical properties in order to achieve superior sensitivities, such as nitrogen–vacancy-centre-based magnetometry at the level of single neurons. By their nature, these wearables typically generate time series data and could thus benefit from such quantum (machine learning) methods.
@@ 3.2.2 Omics
Omics time series data is particularly valuable for biomarker discovery, as it allows one to observe temporal patterns and gene-gene/protein-protein interactions. Examples includes disease progression/sub-typingand patient response to treatments.
Non-invasive approaches to use liquid biopsies allow one to capture a mixture of the omics profile of cancer cells in a patient via cell-free DNA (cfDNA). Recently, studies have demonstrated the application of cfDNA to identify drug resistance mechanisms [72]. Moreover, in the last few years several studies have shown that cfDNA fragmentomics’ characteristics differ in normal and diseased individuals without the need to distinguish the source of the cfDNA fragments, which makes it a promising novel biomarker in particular for early cancer detection [73, 74]. Indeed, ML algorithms have been applied to such tasks with promising results to identify late-stage cancer [73] but still fail to detect early signs in the development of the disease. Another recent application of omics time series is in the case of epilepsy post-surgical seizure recurrence [75] where omics data were used to subtype patients based on their treatment responses.
In general, classical ML often cannot be easily applied to omics-based time series data due to issues such as small sample numbers or missing data. Therefore, data imputation is required, and current imputation techniques often do not incorporate time-dependence, thereby leaving out critical information during imputation. Recent work has demonstrated that graph networks that model topology across samples can be hugely beneficial in such cases [76]. An example is the use of dynamic Bayesian networks for analysis of time series microbiome data, which has been demonstrated to be a critical biomarker for many diseases states [77, 78]. Still, these topological methods quickly become computationally infeasible with increasingly high-dimensional and complicated data sets, which is where QML and quantum topological data analysis (QTDA)approaches may be able to help, providing an orthogonal point of view and insights. Indeed, they could be used to differentiate with high accuracy between healthy individuals and cancer patients in early stages of diseases, leading to more targeted and effective treatments.
@@ 3.2.3 Medical images
QRC techniques are relevant techniques for application to these problems. However, the quantum computers would have to be larger than the current generations or the initial image data would have to be downscaled with classical pre-processing techniques, such as passing the image through a residual networkbefore sending it through the QRC model. Further investigation is required to understand whether the loss of information during the classical pre-processing would negate the potential advantages of using QRC and related quantum methods instead of classical techniques.
@@ 3.3 Erroneous data
Noise and errors often appear in the context of quantum computing due to the fragility of the quantum systems and their inherent tendency to decohere. A variety of approaches is being researched in order to handle, or even leverage, such noise[79] and progress towards quantum error correction, arguably the holy grail of quantum computing. The focus of this discussion, however, is on handling noisy datasets where the labels are incorrect or missing, which represents a major challenge for classical algorithms. Such misannotations are also of concern when data is synthetically generated [80], an area that is also being explored with QGANs and related methods.
@@ 3.3.1 EHRs
EHR data has been notoriously difficult to analyze due to factors such as errors, lack of interoperability, record gaps, and unusually large or small data sizes. While leveraging such data may also allow better generalization behavior through classical modelsQML has shown promise in dealing with such difficult datasets even more effectively. For example, the usual presence of quantum noise may actually lead to a greater degree of global robustness and thus fairer models [81]. Moreover, quantum transfer learning may be beneficial when the (real-world) dataset of interest is particularly low- or high-dimensional [82]. Quantum support vector machines have been applied to EHRs and have, for small datasets, been shown to be competitive with classical approaches in classifying ischemic heart disease [83] and predicting the persistence of rheumatoid arthritis patients [84]. Small datasets, for instance for rare-disease or other very restricted cohorts, may in fact lead to some of the earliest quantum advantages, given that quantum generative and other QML models are particularly suitable for such a setting [17, 85]. Finally, the process of imputation itself may be enhanced, leveraging quantum determinantal point processes [86].
@@ 4 Open research challenges
In the space of biomarker discovery, the loading of data is one key challenge. As described before, biomarker discovery often necessitates extracting insights from large classical datasets. Despite progress with QRAMs, breakthroughs in this area will unlock many more applications. Related to that, quantum algorithms remain an intensely researched area. New ones are being continually discovered; for instance, a study recently claimed an exponential advantage in pathfinding for graphs [93]. Furthermore, quantum federated learning represents another novel area [94, 95, 96, 97]. The idea is to enhance federated learning models where different entities train a model while keeping their data decentralized. Given the growing importance federated learning has for healthcare and medicine, from allowing multi-institutional collaborations without patient data sharingto enabling digital healthto enhancing the internet of medical things, this is a clear innovation opportunity.
In the area of biomarker validation, a critical aspect concerns benchmarks. The challenge with demonstrating value from quantum computing is that the (classical) goal posts are often not clearly defined and may shift (particularly in response to, or even inspired by, quantum algorithm advances). This makes it challenging to justify the efforts required to develop and use quantum algorithm-based tools. Connected with that, generalization is a key issue due to individual biomarker variability resulting from genetic and lifestyle factors. As a result, more granular (quantum) models will have to be developed, which might again require more (classical) data and novel quantum algorithms.
@@ 5 Conclusion
Clearly, there are also still discrepancies with regard to the access to advanced forms of computing, such as quantum computing, across the world [98]. While quantum computing, in particular, has been rapidly made more widely available, it is likely to remain relatively expensive for some time to come; today, typical costs are thousands of US dollars per QPU hour. Thus, democratization of such transformative technologies is an important issue that also needs to be considered in order to realize their full impact across the world.
6 Acknowledgments
Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525. This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.